# InfinityStar: Unified Spacetime AutoRegressive Modeling for Visual Generation

### v1

@@ 1 Introduction
Visual synthesis has witnessed remarkable progress in recent years, largely propelled by the scaling of Transformer architectures. In particular, video generation has attracted growing interest from both academia and industry, owing to its wide-ranging applications in content creation, world simulation, etc. At present, diffusion models[3, 20, 19, 32, 9, 47] lead the field by iteratively denoising latent representations to produce high-fidelity clips. Concurrently, autoregressive models[18, 34, 10] have been explored for their potential to unify image and video generation and to generalize over longer time horizons.
@@
These observations motivate the need for a generation framework that simultaneously possess high visual quality, efficiency and temporal generalization. Recently, Visual AutoRegressive modeling (VAR)[29] redefined image generation as a coarse-to-fine next-scale prediction. Its follow-up work, Infinity [15] further introduces bitwise modeling and scales up the vocabulary size, achieving comparable performance to diffusion models while offering significant advantages in inference speed. Inspired by the success of VAR [29] and Infinity [15], we present InfinityStar, a Spacetime Pyramid Modeling for unified text‐to‐image, text‐to‐video, zero-shot image‐to‐video, and zero‐shot video extrapolation. This framework models a video as an image pyramid and multiple clip pyramids, not only naturally inheriting the text-to-image capabilities but also decoupling static appearance from dynamic motions in videos. Furthermore, we introduce several key improvements. First, we improve discrete reconstruction quality by leveraging knowledge inheritance from a continuous video tokenizer. Second, we introduce Stochastic Quantizer Depth during training of the tokenizer to alleviate the imbalanced information distribution across scales. Third, we propose Semantic Scales Repetition, which refines the predictions of early semantic scales in a video, significantly enhancing fine-grained details and complex motions of the generated videos.
We train InfinityStar on large‐scale video corpora to support generating videos of up to 720p resolution and variable durations. On the VBench benchmark[45], InfinityStar establishes a new state-of-the-art among autoregressive video generation models, even surpassing industry-leading HunyuanVideo[19] (83.74 v.s 83.24). Besides, InfinityStar shows a great advantage in terms of speed. Using visual tokenizers of the same compression rate, InfinityStar achieves a $10\times$ reduction in inference latency relative to leading diffusion models.
@@ 2.1 Video Diffusion Models
Diffusion models excel at generating high-fidelity data by gradually denoising random noise and has been widely applied in video generation. Early attempts [2, 4, 43] are built on U-Net architectures, demonstrating the feasibility of this approach but falling short in producing sharp, temporally coherent frames due to limited model capacity. The advent of Diffusion Transformers (DiT [23]) marked a turning point. SORA [3] harnessed DiT’s scaling ability to process spatio-temporal patches at scale, dramatically improving both video consistency and generation quality. The success of SORA has catalyzed a wave of innovation [40, 19, 32, 47] across the industry, propelling video generation to unprecedented levels of coherence and fidelity. Although video diffusion models deliver outstanding quality, their slow generation speed hinders the production of high-resolution, long-duration videos.
@@ 2.2 Video AutoRegressive Models
Another class of methods [34, 10, 18] employs autoregressive models for video generation. Inspired by the success of LLMs, these works predict video tokens in specific orders using an autoregressive Transformer. For example, Emu3 [34] performs next-token prediction along both spatial and temporal axes, while Nova [10] first predicts spatial tokens set-by-set and subsequently proceeds frame-by-frame in the temporal dimension. Although achieving preliminary progress, they require hundreds to thousands of inference steps, resulting in prohibitively low generation efficiency. In contrast, recent advances in next-scale prediction [29, 15] have demonstrated state-of-the-art performance in image synthesis, offering both improved quality and markedly faster inference. In this work, we extend the next-scale prediction paradigm to the unified image and video generation.
@@ 2.3 Discrete Video Tokenizers
For a long time, discrete [39, 41] and continuous [19, 40, 32] video tokenizers have been developed independently. Although some works [1, 33] provide both discrete and continuous tokenizers, the network configurations are usually not aligned. For example, Cosmos [1] chooses 6 and 16 as latent dimensions in its discrete and continuous variants respectively. This misalignment hinders the knowledge reuse between two types of tokenizers. As a result, most mainstream discrete video tokenizers are either trained from scratch [1] or starting from a pretrained discrete image tokenizer [41, 33]. However, these training strategies have the following drawbacks. First, training from scratch is inefficient and converges slowly. Second, weights pretrained on static images are not optimal for video reconstruction. To alleviate these deficiencies, we propose a new training strategy, which inherits the architecture and knowledge of a trained continuous video tokenizer. Experiments show that this strategy significantly boosts the convergence of discrete video tokenizers.
@@ 3.1 Preliminaries
Infinity for Image Generation. Infinity [15] decomposes an image into a sequence of hierarchical token blocks using a visual tokenizer and models the relationship between tokens by a visual autoregressive Transformer (VAR Transformer). To cover images of various sizes, Infinity pre-defines a list of token block sizes $\{(h_{1},w_{1}),...(h_{K},w_{K})\}$, called scale schedule. The size $(h_{i},w_{i})$ in scale schedule grows as $i$ increases, forming a pyramid-like structure, which we refer as image pyramid in later discussion. Next we introduce the training and inference procedure of Infinity.
In the first training stage, a visual tokenizer learns to reconstruct the raw image and compress it into a sequence of discrete tokens, which can be modeled by the VAR Transformer in the next stage. Specifically, the tokenizer first encodes the raw images into compact latents, then transforms latents into $K$ discrete residual token blocks $(r_{1},r_{2},...,r_{K})$ using a bitwise multi-scale residual quantizer [15]. Each token block $r_{i}$ consists of $h_{i}\times w_{i}$ discrete tokens of $d$-dim with vocabulary size of $2^{d}$. Then in the second stage, a VAR Transformer is trained to predict next residual token block $r_{k}$ conditioned on text embedding $\psi(t)$ and former tokens blocks $r_{<k}$. Formally, in each step, VAR Transformer predicts a conditional probability $p(r_{k}|r_{<k},\psi(t))$. During the inference, Infinity generates an image by running the VAR Transformer $K$ times autoregressively, merging the predicted tokens and running the tokenizer decoder once.
@@ 3.2 Spacetime Pyramid Modeling for Unified Generation
Extending the spatial-only next-scale prediction paradigm of Infinity [15] to video generation presents a primary challenge: how to incorporate the temporal dimension. The straightforward strategies are either letting time grows uniformly, i.e., from $(1,1,1)$ to $(T,H,W)$, or keeping time constant, i.e., from $(T,1,1)$ to $(T,H,W)$. We empirically found that letting time grow uniformly produces flickering videos. As for the constant time pyramid, we refer to it as the pseudo-spacetime pyramid. Despite its conceptual simplicity, it suffers from two fundamental limitations. First, the treatment of videos differs markedly from that of images, preventing a text-to-video (T2V) model from effectively leveraging the knowledge learned by a text-to-image (T2I) model and complicating its extension to tasks such as image-to-video (I2V). Second, because appearance and motion in videos are coupled in this design, the model faces significant difficulty in accurately fitting both aspects.
To overcome these challenges, we propose a novel spacetime pyramid modeling framework as shown in Fig.1. Each video is decomposed into sequential clips $\{c_{1},c_{2},\cdots,c_{N}\}$. We regard the first frame as $c_{1}$ (i.e., $T=1$) to specifically encode static appearance cues in videos and other clips share an equal duration $T>1$. Each clip is modeled as a 3D volume pyramid similar to Infinity [15]. In particular, for each clip, there are $K$ scales with each represented as a residual token block $r_{k}$ of $(T,h_{k},w_{k})$ dimension. It is worth noting that all scales in the pyramid are extended only in spatial dimension instead of time. Mathematically, the tokens in the first clip are generated auto-regressively across scales as:
@@ 3.3 Visual Tokenizer
Knowledge Inheritance from Continuous Video Tokenizer. Instead of designing and training a discrete video tokenizer from scratch, we inherit the architecture and weights of a trained continuous video tokenizer, i.e. video VAE. Specifically, we first insert a parameter-free quantizer between the pre-trained VAE encoder and the decoder. The quantizer is based on binary spherical quantization [49], being similar to that of Infinity [15] but with new spacetime pyramid scale schedule. This does not introduce any new parameter like codebook in VQ [30] and well retains knowledge of the original VAE. As shown in Fig.2, the discrete video tokenizer reconstructs videos decently, even without any fine-tuning. To further improve the reconstruction quality, we fine-tune the new tokenizer jointly on images and videos like previous works [33, 1]. During the fine-tuning, the KL loss of the original VAE is replaced with the commitment loss plus the entropy penalty [49]. As shown in Fig.2, with the help of knowledge of continuous video VAE, the convergence accelerates dramatically.
@@ 3.4 Spacetime Autoregressive Transformer
Spacetime Sparse Attention. Autoregressive video generation faces significant challenges due to the high computational costs of long context. As on the left of Fig.4, Infinity [15] employs a block-wise causal mask for single pyramid modeling. Switti [31] verifies that conditioning solely on inputs from preceding scales is sufficient for next-scale predictions, resulting in a sparser attention mask as on the middle of Fig.4. For long video generation, it’s necessary to attend history tokens to achieve temporal consistency. However, attending full history leads to an explosively long sequence. Considering each clip corresponds to 5s, which is sufficient to maintain temporal consistency, here we only attend to the last scale of the preceding clip. Finally, we obtain a highly sparse attention as show in Fig.4 (right). Our spacetime sparse attention drastically reduces computational overhead in attention during both training and inference, while delivering better performance.
@@ 4.2 Text-to-Image Generation
The upper part of Fig.5 shows images generated by our InfinityStar-T2I model, showcasing InfinityStar’s strength in generating high-fidelity and photo-realistic images across various categories and styles. We also carry out the quantitative evaluation on the GenEval[14] and DPG[16] benchmarks. As in Tab.1, InfinityStar achieves the best overall score of 0.79 on the GenEval bench with a prompt rewriter. It’s worth noting that InfinityStar exceeds Infinity by 6% on overall score. We attribute the significant improvement to the larger model size and the architectural innovations. On the DPG bench, InfinityStar reaches an overall score of 86.55, surpassing Infinity by 3.09%. These quantitative results demonstrate InfinityStar’s strong capabilities of image generation following users’ prompts.
Table 1: Evaluation on the GenEval [14] and DPG [16] benchmark. $\dagger$ result is with prompt rewriting or self-CoT.
@@
SDXL [24] 2.6B 0.74 0.15 0.23 0.55 83.27 86.76 74.7
PixArt-Sigma [5] 0.6B 0.62 0.14 0.27 0.55 86.89 86.59 80.5
SD3 (d=38) [11] 8B 0.89 0.34 0.47 0.71 - - -
Goku [6] 2B - - - 0.76 - - 83.6
Transfusion [51] 7.3B - - - 0.63 - - -
SANA-1.0 [37] 1.6B - - - 0.66 - - 84.8
FLUX-dev [21] 12B - - - 0.67 - - 84.0
FLUX-schnell [21] 12B - - - 0.71 - - 84.8
@@
LlamaGen [26] 0.8B 0.34 0.07 0.04 0.32 65.2
Chameleon [27] 7B - - - 0.39 - - -
Show-o [38] 1.3B 0.80 0.31 0.50 0.68 - - 67.5
Liquid [36] 7B 0.73 0.17 0.37 0.55 - - -
UniTok [22] 7B 0.71 0.26 0.45 0.59 - - -
Janus [35] 1.3B 0.68 0.46 0.42 0.61 - - -
Emu3 [34] 8B 0.81 0.49 0.45 0.66 - - 81.6
Fluid [12] 10.5B 0.83 0.39 0.51 0.69 - - -
NextStep-1 [28] 14B - - - 0.73 - - 85.28
Infinity [15] 2B 0.85 0.49 0.57 0.73 93.11 90.76 83.46
@@
VideoCrafter-2.0[4] 1.5B 95.00 55.29 40.66 25.13 82.20 73.42 80.44
OpenSora V1.2[50] 1.1B 85.80 42.47 58.41 23.89 80.71 73.30 79.23
Show-1[44] 6B 95.60 47.03 45.47 23.06 80.42 72.98 78.93
Gen-3 [13] - 96.40 54.57 53.64 24.31 84.11 75.17 82.32
CogVideoX-5B[40] 5B 99.40 53.20 62.11 24.91 82.75 77.04 81.61
HunyuanVideo[19] 13B 94.40 53.88 68.55 19.80 85.09 75.82 83.24
Goku[6] 2B 97.60 57.08 79.48 23.08 85.60 81.87 84.85
Wan 2.1[32] 14B 98.80 53.67 81.44 21.13 85.64 80.95 84.70

### v2

@@ 1 Introduction
Visual synthesis has witnessed remarkable progress in recent years, largely propelled by the scaling of Transformer architectures. In particular, video generation has attracted growing interest from both academia and industry, owing to its wide-ranging applications in content creation, world simulation, etc. At present, diffusion models[sora, kling, Hunyuanvideo, Wan, veo3, waver] lead the field by iteratively denoising latent representations to produce high-fidelity clips. Concurrently, autoregressive models[videopoet, wang2024emu3, nova] have been explored for their potential to unify image and video generation and to generalize over longer time horizons.
@@
These observations motivate the need for a generation framework that simultaneously possess high visual quality, efficiency and temporal generalization. Recently, Visual AutoRegressive modeling (VAR)[VAR] redefined image generation as a coarse-to-fine next-scale prediction. Its follow-up work, Infinity [Infinity] further introduces bitwise modeling and scales up the vocabulary size, achieving comparable performance to diffusion models while offering significant advantages in inference speed. Inspired by the success of VAR [VAR] and Infinity [Infinity], we present InfinityStar, a Spacetime Pyramid Modeling for unified text‐to‐image, text‐to‐video, zero-shot image‐to‐video, and zero‐shot video extrapolation. This framework models a video as an image pyramid and multiple clip pyramids, not only naturally inheriting the text-to-image capabilities but also decoupling static appearance from dynamic motions in videos. Furthermore, we introduce several key improvements. First, we improve discrete reconstruction quality by leveraging knowledge inheritance from a continuous video tokenizer. Second, we introduce Stochastic Quantizer Depth during training of the tokenizer to alleviate the imbalanced information distribution across scales. Third, we propose Semantic Scales Repetition, which refines the predictions of early semantic scales in a video, significantly enhancing fine-grained details and complex motions of the generated videos.
We train InfinityStar on large‐scale video corpora to support generating videos of up to 720p resolution and variable durations. On the VBench benchmark[vbench], InfinityStar establishes a new state-of-the-art among autoregressive video generation models, even surpassing industry-leading HunyuanVideo[Hunyuanvideo] (83.74 v.s 83.24). Besides, InfinityStar shows a great advantage in terms of speed. Using visual tokenizers of the same compression rate, InfinityStar achieves a $10\times$ reduction in inference latency relative to leading diffusion models.
@@ 2.1 Video Diffusion Models
Diffusion models excel at generating high-fidelity data by gradually denoising random noise and has been widely applied in video generation. Early attempts [svd, videocrafter, pixeldance] are built on U-Net architectures, demonstrating the feasibility of this approach but falling short in producing sharp, temporally coherent frames due to limited model capacity. The advent of Diffusion Transformers (DiT [dit]) marked a turning point. SORA [sora] harnessed DiT’s scaling ability to process spatio-temporal patches at scale, dramatically improving both video consistency and generation quality. The success of SORA has catalyzed a wave of innovation [cogvideox, Hunyuanvideo, Wan, waver] across the industry, propelling video generation to unprecedented levels of coherence and fidelity. Although video diffusion models deliver outstanding quality, their slow generation speed hinders the production of high-resolution, long-duration videos.
@@ 2.2 Video AutoRegressive Models
Another class of methods [wang2024emu3, nova, videopoet] employs autoregressive models for video generation. Inspired by the success of LLMs, these works predict video tokens in specific orders using an autoregressive Transformer. For example, Emu3 [wang2024emu3] performs next-token prediction along both spatial and temporal axes, while Nova [nova] first predicts spatial tokens set-by-set and subsequently proceeds frame-by-frame in the temporal dimension. Although achieving preliminary progress, they require hundreds to thousands of inference steps, resulting in prohibitively low generation efficiency. In contrast, recent advances in next-scale prediction [VAR, Infinity] have demonstrated state-of-the-art performance in image synthesis, offering both improved quality and markedly faster inference. In this work, we extend the next-scale prediction paradigm to the unified image and video generation.
@@ 2.3 Discrete Video Tokenizers
For a long time, discrete [videogpt, magvit2] and continuous [Hunyuanvideo, cogvideox, Wan] video tokenizers have been developed independently. Although some works [cosmos, omnitokenizer] provide both discrete and continuous tokenizers, the network configurations are usually not aligned. For example, Cosmos [cosmos] chooses 6 and 16 as latent dimensions in its discrete and continuous variants respectively. This misalignment hinders the knowledge reuse between two types of tokenizers. As a result, most mainstream discrete video tokenizers are either trained from scratch [cosmos] or starting from a pretrained discrete image tokenizer [magvit2, omnitokenizer]. However, these training strategies have the following drawbacks. First, training from scratch is inefficient and converges slowly. Second, weights pretrained on static images are not optimal for video reconstruction. To alleviate these deficiencies, we propose a new training strategy, which inherits the architecture and knowledge of a trained continuous video tokenizer. Experiments show that this strategy significantly boosts the convergence of discrete video tokenizers.
@@ 3.1 Preliminaries
Infinity for Image Generation. Infinity [Infinity] decomposes an image into a sequence of hierarchical token blocks using a visual tokenizer and models the relationship between tokens by a visual autoregressive Transformer (VAR Transformer). To cover images of various sizes, Infinity pre-defines a list of token block sizes $\{(h_{1},w_{1}),...(h_{K},w_{K})\}$, called scale schedule. The size $(h_{i},w_{i})$ in scale schedule grows as $i$ increases, forming a pyramid-like structure, which we refer as image pyramid in later discussion. Next we introduce the training and inference procedure of Infinity.
In the first training stage, a visual tokenizer learns to reconstruct the raw image and compress it into a sequence of discrete tokens, which can be modeled by the VAR Transformer in the next stage. Specifically, the tokenizer first encodes the raw images into compact latents, then transforms latents into $K$ discrete residual token blocks $(r_{1},r_{2},...,r_{K})$ using a bitwise multi-scale residual quantizer [Infinity]. Each token block $r_{i}$ consists of $h_{i}\times w_{i}$ discrete tokens of $d$-dim with vocabulary size of $2^{d}$. Then in the second stage, a VAR Transformer is trained to predict next residual token block $r_{k}$ conditioned on text embedding $\psi(t)$ and former tokens blocks $r_{<k}$. Formally, in each step, VAR Transformer predicts a conditional probability $p(r_{k}|r_{<k},\psi(t))$. During the inference, Infinity generates an image by running the VAR Transformer $K$ times autoregressively, merging the predicted tokens and running the tokenizer decoder once.
@@ 3.2 Spacetime Pyramid Modeling for Unified Generation
Extending the spatial-only next-scale prediction paradigm of Infinity [Infinity] to video generation presents a primary challenge: how to incorporate the temporal dimension. The straightforward strategies are either letting time grows uniformly, i.e., from $(1,1,1)$ to $(T,H,W)$, or keeping time constant, i.e., from $(T,1,1)$ to $(T,H,W)$. We empirically found that letting time grow uniformly produces flickering videos. As for the constant time pyramid, we refer to it as the pseudo-spacetime pyramid. Despite its conceptual simplicity, it suffers from two fundamental limitations. First, the treatment of videos differs markedly from that of images, preventing a text-to-video (T2V) model from effectively leveraging the knowledge learned by a text-to-image (T2I) model and complicating its extension to tasks such as image-to-video (I2V). Second, because appearance and motion in videos are coupled in this design, the model faces significant difficulty in accurately fitting both aspects.
To overcome these challenges, we propose a novel spacetime pyramid modeling framework as shown in Fig.1. Each video is decomposed into sequential clips $\{c_{1},c_{2},\cdots,c_{N}\}$. We regard the first frame as $c_{1}$ (i.e., $T=1$) to specifically encode static appearance cues in videos and other clips share an equal duration $T>1$. Each clip is modeled as a 3D volume pyramid similar to Infinity [Infinity]. In particular, for each clip, there are $K$ scales with each represented as a residual token block $r_{k}$ of $(T,h_{k},w_{k})$ dimension. It is worth noting that all scales in the pyramid are extended only in spatial dimension instead of time. Mathematically, the tokens in the first clip are generated auto-regressively across scales as:
@@ 3.3 Visual Tokenizer
Knowledge Inheritance from Continuous Video Tokenizer. Instead of designing and training a discrete video tokenizer from scratch, we inherit the architecture and weights of a trained continuous video tokenizer, i.e. video VAE. Specifically, we first insert a parameter-free quantizer between the pre-trained VAE encoder and the decoder. The quantizer is based on binary spherical quantization [BSQ], being similar to that of Infinity [Infinity] but with new spacetime pyramid scale schedule. This does not introduce any new parameter like codebook in VQ [vqvae] and well retains knowledge of the original VAE. As shown in Fig.2, the discrete video tokenizer reconstructs videos decently, even without any fine-tuning. To further improve the reconstruction quality, we fine-tune the new tokenizer jointly on images and videos like previous works [omnitokenizer, cosmos]. During the fine-tuning, the KL loss of the original VAE is replaced with the commitment loss plus the entropy penalty [BSQ]. As shown in Fig.2, with the help of knowledge of continuous video VAE, the convergence accelerates dramatically.
@@ 3.4 Spacetime Autoregressive Transformer
Spacetime Sparse Attention. Autoregressive video generation faces significant challenges due to the high computational costs of long context. As on the left of Fig.4, Infinity [Infinity] employs a block-wise causal mask for single pyramid modeling. Switti [Switti] verifies that conditioning solely on inputs from preceding scales is sufficient for next-scale predictions, resulting in a sparser attention mask as on the middle of Fig.4. For long video generation, it’s necessary to attend history tokens to achieve temporal consistency. However, attending full history leads to an explosively long sequence. Considering each clip corresponds to 5s, which is sufficient to maintain temporal consistency, here we only attend to the last scale of the preceding clip. Finally, we obtain a highly sparse attention as show in Fig.4 (right). Our spacetime sparse attention drastically reduces computational overhead in attention during both training and inference, while delivering better performance.
@@ 4.2 Text-to-Image Generation
The upper part of Fig.5 shows images generated by our InfinityStar-T2I model, showcasing InfinityStar’s strength in generating high-fidelity and photo-realistic images across various categories and styles. We also carry out the quantitative evaluation on the GenEval[ghosh2024geneval] and DPG[DPG-bench] benchmarks. As in Tab.1, InfinityStar achieves the best overall score of 0.79 on the GenEval bench with a prompt rewriter. It’s worth noting that InfinityStar exceeds Infinity by 6% on overall score. We attribute the significant improvement to the larger model size and the architectural innovations. On the DPG bench, InfinityStar reaches an overall score of 86.55, surpassing Infinity by 3.09%. These quantitative results demonstrate InfinityStar’s strong capabilities of image generation following users’ prompts.
Table 1: Evaluation on the GenEval [ghosh2024geneval] and DPG [DPG-bench] benchmark. $\dagger$ result is with prompt rewriting or self-CoT.
@@
SDXL [sdxl] 2.6B 0.74 0.15 0.23 0.55 83.27 86.76 74.7
PixArt-Sigma [chen2024pixart_sigma] 0.6B 0.62 0.14 0.27 0.55 86.89 86.59 80.5
SD3 (d=38) [stable-diffusion3] 8B 0.89 0.34 0.47 0.71 - - -
Goku [goku] 2B - - - 0.76 - - 83.6
Transfusion [transfusion] 7.3B - - - 0.63 - - -
SANA-1.0 [xie2024sana] 1.6B - - - 0.66 - - 84.8
FLUX-dev [FLUX] 12B - - - 0.67 - - 84.0
FLUX-schnell [FLUX] 12B - - - 0.71 - - 84.8
@@
LlamaGen [llamagen] 0.8B 0.34 0.07 0.04 0.32 65.2
Chameleon [chameleon-meta] 7B - - - 0.39 - - -
Show-o [show-o] 1.3B 0.80 0.31 0.50 0.68 - - 67.5
Liquid [liquid] 7B 0.73 0.17 0.37 0.55 - - -
UniTok [unitok] 7B 0.71 0.26 0.45 0.59 - - -
Janus [janus] 1.3B 0.68 0.46 0.42 0.61 - - -
Emu3 [wang2024emu3] 8B 0.81 0.49 0.45 0.66 - - 81.6
Fluid [fluid] 10.5B 0.83 0.39 0.51 0.69 - - -
NextStep-1 [nextstep1] 14B - - - 0.73 - - 85.28
Infinity [Infinity] 2B 0.85 0.49 0.57 0.73 93.11 90.76 83.46
@@
VideoCrafter-2.0[videocrafter] 1.5B 95.00 55.29 40.66 25.13 82.20 73.42 80.44
OpenSora V1.2[opensora] 1.1B 85.80 42.47 58.41 23.89 80.71 73.30 79.23
Show-1[show-1] 6B 95.60 47.03 45.47 23.06 80.42 72.98 78.93
Gen-3 [Gen-3] - 96.40 54.57 53.64 24.31 84.11 75.17 82.32
CogVideoX-5B[cogvideox] 5B 99.40 53.20 62.11 24.91 82.75 77.04 81.61
HunyuanVideo[Hunyuanvideo] 13B 94.40 53.88 68.55 19.80 85.09 75.82 83.24
Goku[goku] 2B 97.60 57.08 79.48 23.08 85.60 81.87 84.85
Wan 2.1[Wan] 14B 98.80 53.67 81.44 21.13 85.64 80.95 84.70